OpenStudy (anonymous):

Input in standard form the equation of the given line the line with m = -3/2 and b = 6

OpenStudy (anonymous):

thank u tho

OpenStudy (anonymous):

Which function below has the lowest y-intercept?

OpenStudy (anonymous):

Kyle started the summer having read 3 books but plans to read 6 books per month over the summer h (x) 3x+4

OpenStudy (texaschic101):

slope(m) = -3/2 y intercept (b) = 6 the equation is : y = -3/2x + 6...however, we need it in standard form : Ax + By = C y = -3/2x + 6 -- multiply the entire equation by 2 to get rid of the fraction 2y = -3x + 12 -- add 3x to both sides 3x + 2y = 12 <== standard form ===================== h(x) = 3x + 4.....this is the same as y = 3x + 4 and the y intercept here is 4 as far as which one has the lowest y intercept....are you comparing it with the first question ? The reason I ask, is because the first question is not a function...until you turn it to one. Or is there another function you are talking about that is not listed ?

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.